OpenStudy (marcelie):

to integrate sin(2x) do i have to change it to 1/2-cos92x) ?

OpenStudy (danjs):

\[\int\limits \sin(2x)dx\] you dont have to change it , you can do a substitution with u let u=2x du=2dx

OpenStudy (danjs):

so you get sin(u) and for the dx, you get dx=du/2 \[\large \int\limits \sin(u)(\frac{ 1 }{ 2 })du=\frac{ 1 }{ 2 }\int\limits \sin(u)du\]

OpenStudy (danjs):

right?

OpenStudy (marcelie):

okay ill give u the full problem lol this is what i did by using the table

OpenStudy (marcelie):